An average tire used by a transportation company lasts 20,000 km with a standard deviation

of 100 km. Assuming that the distance traveled by a tire is normally distributed, what is the

probability that a tire used by the company will last at most 30,000 km?


1
Expert's answer
2022-04-21T10:37:22-0400

"P(X\le30000)=P(X<30000)=P(Z<\frac{30000-20000}{100})=P(Z<100)=1."